Prove that SL(n,F) is a subgroup of GL(n,F)


1
Expert's answer
2021-06-25T12:27:49-0400

Let us prove that "SL(n,F)=\{A\in GL(n,F):\det A =1\}" is a subgroup of "GL(n,F)." If "A,B\in SL(n,F)," then "\det A=\det B =1." It follows that "\det(AB)=\det A\cdot \det B=1\cdot 1=1," and hence "AB\in SL(n,F)." It follows also that "\det (A^{-1})=\frac{1}{\det A}=\frac{1}{1}=1," and hence "A^{-1}\in SL(n, F)." We conclude that "SL(n,F)"

is a subgroup of "GL(n,F)."